# ATOD: An Evaluation Framework and Benchmark for Agentic Task-Oriented Dialogue System

### v1

@@ Abstract
Recent advances in task-oriented dialogue (TOD) systems, driven by large language models (LLMs) with extensive API and tool integration, have enabled conversational agents to coordinate interleaved goals, maintain long-horizon context, and act proactively through asynchronous execution. These capabilities extend beyond traditional TOD systems, yet existing benchmarks lack systematic support for evaluating such agentic behaviors. To address this gap, we introduce ATOD, a benchmark and synthetic dialogue generation pipeline that produces richly annotated conversations requiring long-term reasoning. ATOD captures key characteristics of Advanced TOD, including multi-goal coordination, dependency management, memory, adaptability, and proactivity. Building on ATOD, we propose ATOD-Eval, a holistic evaluation framework that translates these dimensions into fine-grained metrics and supports reproducible offline and online evaluation. We further present a strong agentic memory-based evaluator for benchmarking on ATOD. Experiments show that ATOD-Eval enables comprehensive assessment across task completion, agentic capability, and response quality, and that the proposed evaluator offers a better accuracy–efficiency trade-off compared to existing memory- and LLM-based approaches under this evaluation setting.
ATOD: An Evaluation Framework and Benchmark for Agentic Task-Oriented Dialogue System
@@ 1 Introduction
To fill this gap, we introduce ATOD, a benchmark and synthetic dialogue generation pipeline that produces richly annotated dialogues requiring long-term recall, interleaved workflows, and explicit goal dependencies. Building on ATOD, we propose ATOD-Eval, a holistic evaluation framework that provides standardized benchmarks and fine-grained metrics for systematically capturing advanced TOD capabilities. ATOD-Eval unifies evaluation and benchmarking by jointly assessing goal completion, dependency management, memory consistency, adaptability, proactivity, and multi-goal coordination, translating these dimensions into reproducible metrics for both offline and online settings. We further present a proposed agentic memory-based evaluator for benchmarking on ATOD, enabling empirical comparison against strong memory- and LLM-based baselines. Extensive experiments validate the proposed benchmark and evaluation framework, showing that the resulting metrics provide comprehensive and consistent assessment of advanced TOD capabilities, while the proposed evaluator consistently outperforms competitive baselines under this evaluation setting.
@@ 3.2 Task Formulation
Formally, let $\mathcal{D}=\{(\mathcal{G}_{i},\mathcal{C}_{i})\}_{i=1}^{N}$ denote a dialogue corpus, where $\mathcal{C}_{i}=\{c_{i,t}\}_{t=1}^{T_{i}}$ is the ordered sequence of dialogue turns and $\mathcal{G}_{i}$ is the associated set of user goals with explicit dependencies. Unlike traditional TOD settings where a goal is confined to a contiguous span, in advanced TOD a single goal $g\in\mathcal{G}_{i}$ may span disjoint intervals of $\mathcal{C}_{i}$, being initiated, suspended, and resumed as the dialogue evolves. We represent each goal with a goal status trajectory $\{\mathrm{Status}(g,t)\}_{t=1}^{T_{i}}$ (e.g., Open $\to$ Pending $\to$ Completed/Failed), capturing a non-contiguous and interleaved lifecycle over extended interactions. The evaluation objective is to assess how well a system manages interdependent goals, maintains long-horizon trajectories, coordinates asynchronous workflows, and provides proactive support in both offline and online settings.
@@ 4.5 Dataset Coverage
ATOD spans diverse domains and goal complexities, ranging from simple two-goal cases to interdependent, long-horizon workflows. Table 2 compares ATOD with existing benchmarks. While prior datasets capture individual aspects of advanced TOD, ATOD uniquely combines multi-goal concurrency, interleaving with asynchronous execution, explicit dependency modeling, proactive behaviors, and turn-level status annotation. This makes ATOD the first dataset purpose-built to comprehensively support evaluation of advanced TOD systems.
@@ 8.3 Metric Validity Analysis
We first summarize the average values of our proposed metrics (Table 4), which reflect different dimensions of system behavior: efficiency, memory, proactivity, and interaction quality. As expected, medium dialogues are shorter and yield higher memory recall, whereas complex dialogues require more turns and show reduced recall, consistent with their higher difficulty.
@@ A.2 Goal Extraction, Co-occurrence Graph Statistics, and Sampling Strategy
We then construct a goal co-occurrence graph where each node is a unique goal and edges represent co-occurrence within the same sequence. Table 8 shows its statistics. The graph contains 52 nodes and 396 edges, forming a single connected component with relatively high density ($0.2986$) and average degree ($15.23$), indicating frequent goal co-occurrence across dialogues. This structure supports diverse sampling of multi-goal trajectories, including high-degree hubs (up to 29) and rare goals (degree as low as 2).
@@ A.3 Complexity Criteria
For categorization process, we follow a three-step procedure. First, goal sampling draws trajectories under a two-category distribution (default: 65% medium, 35% complex). Second, annotation enriches sampled goals with slots, dependencies, and realistic characteristics. Third, hybrid classification assigns complexity using pre-defined rules combined with LLM analysis, considering quantitative factors (goal count, domain diversity, dependency structures), qualitative factors (goal interdependence and coordination complexity), and realistic dialogue requirements such as interleaving and proactivity needs.
@@ A.5 ATOD: Quality Control
We use the following LLM-based quality control prompt to verify goal clarity, slot validity, and annotation consistency of annotated goals (§4.2) prior to dialogue generation.

### v2

@@ Abstract
Recent advances in task-oriented dialogue (TOD) systems, driven by large language models (LLMs) with extensive API and tool integration, have enabled conversational agents to coordinate interleaved goals, maintain long-horizon context, and act proactively through asynchronous execution. These capabilities extend beyond traditional TOD systems, yet existing benchmarks lack systematic support for evaluating such agentic behaviors. To address this gap, we introduce ATOD, a benchmark and synthetic dialogue generation pipeline that produces richly annotated conversations requiring long-term reasoning. ATOD captures key characteristics of Advanced TOD, including multi-goal coordination, dependency management, memory, adaptability, and proactivity. Building on ATOD, we propose ATOD-Eval, a holistic evaluation framework that translates these dimensions into fine-grained metrics and supports reproducible offline and online evaluation. We further propose a strong agentic memory-based evaluator for benchmarking on ATOD. Experiments show that ATOD-Eval enables comprehensive assessment across task completion, agentic capability, and response quality, and that the proposed evaluator offers a better accuracy–efficiency trade-off compared to existing memory- and LLM-based approaches under this evaluation setting.
ATOD: An Evaluation Framework and Benchmark for Agentic Task-Oriented Dialogue Systems
@@ 1 Introduction
To fill this gap, we introduce ATOD, a benchmark and synthetic dialogue generation pipeline that produces richly annotated dialogues requiring long-term recall, interleaved workflows, and explicit goal dependencies. Building on ATOD, we propose ATOD-Eval, a holistic evaluation framework that provides standardized benchmarks and fine-grained metrics for systematically capturing advanced TOD capabilities. ATOD-Eval unifies evaluation and benchmarking by jointly assessing goal completion, dependency management, memory consistency, adaptability, proactivity, and multi-goal coordination, translating these dimensions into reproducible metrics for both offline and online settings. We further present a proposed agentic memory-based evaluator for benchmarking on ATOD, enabling empirical comparison against strong memory- and LLM-based baselines. Extensive experiments validate the proposed benchmark and evaluation framework, showing that the resulting metrics provide a comprehensive and consistent assessment of advanced TOD capabilities, while the proposed evaluator consistently outperforms competitive baselines under this evaluation setting.
@@ 3.2 Task Formulation
Formally, let $\mathcal{D}=\{(\mathcal{G}_{i},\mathcal{C}_{i})\}_{i=1}^{N}$ denote a dialogue corpus, where $\mathcal{C}_{i}=\{c_{i,t}\}_{t=1}^{T_{i}}$ is the ordered sequence of dialogue turns and $\mathcal{G}_{i}$ is the associated set of user goals with explicit dependencies. Unlike traditional TOD settings where a goal is confined to a contiguous span, in advanced TOD, a single goal $g\in\mathcal{G}_{i}$ may span disjoint intervals of $\mathcal{C}_{i}$, being initiated, suspended, and resumed as the dialogue evolves. We represent each goal with a goal status trajectory $\{\mathrm{Status}(g,t)\}_{t=1}^{T_{i}}$ (e.g., Open $\to$ Pending $\to$ Completed/Failed), capturing a non-contiguous and interleaved lifecycle over extended interactions. The evaluation objective is to assess how well a system manages interdependent goals, maintains long-horizon trajectories, coordinates asynchronous workflows, and provides proactive support in both offline and online settings.
@@ 4.5 Dataset Coverage
ATOD spans diverse domains and goal complexities, ranging from simple two-goal cases to interdependent, long-horizon workflows. Table 2 compares ATOD with existing benchmarks. While prior datasets capture individual aspects of advanced TOD, ATOD uniquely combines multi-goal concurrency, interleaving with asynchronous execution, explicit dependency modeling, proactive behaviors, and turn-level status annotation. This makes ATOD the first dataset purpose-built to comprehensively support the evaluation of advanced TOD systems.
@@ 8.3 Metric Validity Analysis
We first summarize the average values of our proposed metrics (Table 4), which reflect different dimensions of system behavior: efficiency, memory, proactivity, and interaction quality. As expected, medium dialogues are shorter and yield higher memory recall, whereas complex dialogues require more turns and show reduced recall, consistent with their greater difficulty.
@@ A.2 Goal Extraction, Co-occurrence Graph Statistics, and Sampling Strategy
We then construct a goal co-occurrence graph where each node is a unique goal and edges represent co-occurrence within the same sequence. Table 8 shows its statistics. The graph contains 52 nodes and 396 edges, forming a single connected component with relatively high density ($0.2986$) and average degree ($15.23$), indicating frequent goal co-occurrence across dialogues. This structure supports a diverse sampling of multi-goal trajectories, including high-degree hubs (up to 29) and rare goals (degree as low as 2).
@@ A.3 Complexity Criteria
For the categorization process, we follow a three-step procedure. First, goal sampling draws trajectories under a two-category distribution (default: 65% medium, 35% complex). Second, annotation enriches sampled goals with slots, dependencies, and realistic characteristics. Third, hybrid classification assigns complexity using pre-defined rules combined with LLM analysis, considering quantitative factors (goal count, domain diversity, dependency structures), qualitative factors (goal interdependence and coordination complexity), and realistic dialogue requirements such as interleaving and proactivity needs.
@@ A.5 ATOD: Quality Control
We use the following LLM-based quality control prompt to verify goal clarity, slot validity, and annotation consistency of annotated goals (§4.2) before dialogue generation.